### v3

# Exact computation of Transfer Entropy with Path Weight Sampling

Abstract
The ability to quantify the directional flow of information is vital to understanding natural systems and designing engineered information-processing systems. A widely used measure to quantify this information flow is the transfer entropy. However, until now, this quantity could only be obtained in dynamical models using approximations that are typically uncontrolled. Here we introduce a computational algorithm called Transfer Entropy-Path Weight Sampling (TE-PWS), which makes it possible, for the first time, to quantify the transfer entropy and its variants exactly for any stochastic model, including those with multiple hidden variables, nonlinearity, transient conditions, and feedback. By leveraging techniques from polymer and path sampling, TE-PWS efficiently computes the transfer entropy as a Monte-Carlo average over signal trajectory space. We apply TE-PWS to linear and nonlinear systems to reveal how transfer entropy can overcome naive applications of the data processing inequality in the presence of feedback.
preprint: APS/123-QED
Information transfer between noisy signals underlies the functionality of diverse natural and man-made networks such as those in biochemical signaling, neuroscience, ecology, wireless communication and finance. Information theory has so far provided a useful framework for quantifying information transmission. In the presence of feedback loops in the network, information travels in both directions between the input and the output. An information-theoretic measure that can quantify the information transfer separately in either direction is the transfer entropy1, 2. Transfer entropy and its variants, such as directed information, conditional transfer entropy or filtered transfer entropy, have been widely used to gain knowledge about the connectivity of a network3, infer causal relations from experiments4, 5, establish fundamental bounds on network performance6, and estimate the minimal physical work required for a computation7, 8. Hence, for a wide range of problems, it is vital to be able to accurately quantify transfer entropies.
However, there are currently no exact methods to compute the transfer entropy in a general many-variable dynamical model. The transfer entropy depends on probability distributions of signal trajectories. Estimating these distributions by binning the experimental or simulated trajectories in histograms is not feasible, as the dimensionality of the trajectory space scales exponentially with signal duration 9, 10. As a result, different approximations are currently being used: either the full history dependence of the transfer entropy is truncated11, or an arbitrary distance metric in trajectory space is chosen for clustering the trajectories12, or a linear or low-order moment-closure approximation is employed13, 14, 15. These approximations can result in uncontrolled errors in complex, nonlinear, many-variable systems16, 17.
Here we fill this gap by introducing TE-PWS, a numerical algorithm to estimate transfer entropies exactly for any stochastic model, including diffusive and jump processes. The estimate is exact, i.e., it is an unbiased statistical estimate of the transfer entropy. TE-PWS can therefore provide ground truth results for any given model. TE-PWS builds on the recently developed PWS algorithm for computing the mutual information between trajectories18. TE-PWS exploits the idea that path likelihoods can be obtained analytically from the Langevin or master equation, from which the transfer entropy is then computed via Monte-Carlo averaging in trajectory space. Additionally, long trajectories are sampled with an importance sampling scheme19, solving the problem of exponential scaling of the computational cost with trajectory duration. We have applied TE-PWS to compute the transfer entropy in a three-variable motif in the presence of feedback, for both linear and nonlinear systems. We show that TE-PWS reproduces analytical results when available and produces novel insights on how information feedback can amplify information transfer. Specifically, the transfer entropy from an input to an output node can overcome a naive application of the data processing inequality even when the mutual information obeys one.
Diffusive process. We introduce transfer entropy and TE-PWS first in the context of diffusive processes. Consider a $d$-dimensional diffusive process $\mathbf{X}(t)$ modelled as a function of time $t$ by a Langevin equation
$\dot{\mathbf{X}}(t)=\mathbf{F}(t)+\bm{\xi}(t),$ (1)
with $\mathbf{F}(t)$ a general drift, and $\bm{\xi}(t)$ a $d$-dimensional Gaussian white noise with a diffusion constant matrix $\mathbf{D}=[D_{ij}]$ such that $\langle\xi_{i}(t)\xi_{j}(t^{{}^{\prime}})\rangle=2D_{ij}\delta(t-t^{{}^{\prime}})$. The drift may depend on the entire past history as well as on time. The transfer entropy from $X_{i}$ to $X_{j}$ over $N$ timesteps of durations $\delta t$ each is defined as1
$\mathcal{T}_{X_{i}\to X_{j}}=\sum_{k=0}^{N-1}I\left(X_{j}(k+1);X_{i,[0,k]}|X_{j,[0,k]}\right)$ (2)
where the index $k$ goes over individual timesteps, $X_{j}(k)$ denotes $X_{j}$ after $k$ timesteps, $X_{j,[0,k]}$ denotes the trajectory over the first $k$ timesteps, and $I(A;B)$ denotes the mutual information between two random variables $A$ and $B$. $\mathcal{T}_{X_{i}\to X_{j}}$ measures the information transferred from the past trajectory of $X_{i}$ to the new updates of $X_{j}$ at every timestep, given the past trajectory of $X_{j}$ is already known. In case the dynamics relaxes into a steady-state, we will also talk about the transfer entropy rate, $\dot{\mathcal{T}}_{X_{i}\to X_{j}}=\lim_{N\to\infty}\mathcal{T}_{X_{i}\to X_{j}}/(N\delta t)$.
We can rewrite the transfer entropy equivalently as
$\mathcal{T}_{X_{i}\to X_{j}}=$ $\sum_{k}\Delta\mathcal{T}^{i\to j}_{[k,k+1]}=\sum_{k}H\left(X_{j}(k+1)|X_{j,[0,k]}\right)$
$\qquad-H\left(X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]}\right)$ (3)
$=$ $\sum_{k}\Bigg{\langle}\ln\frac{P\left(X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]}\right)}{P\left(X_{j}(k+1)|X_{j,[0,k]}\right)}\Bigg{\rangle}$ (4)
where $\Delta\mathcal{T}^{i\to j}_{[k,k+1]}$ denotes the increment at every timestep, $H(A)$ denotes the Shannon entropy associated with the probability distribution $P(A)$ of $A$, and the angular brackets denote an average over the joint probability $P\left(X_{i,[0,N]},X_{j,[0,N]}\right)$. Eq. 3 shows that transfer entropy quantifies the additional information in $X_{j}(k+1)$ that arrives from $X_{i,[0,k]}$ beyond that which is already present in the past trajectory $X_{j,[0,k]}$. This occurs either through direct causal action, or through a third variable $X_{l}$, schematically demonstrated in Fig. 1a. If $X_{i}$ does not affect the dynamics of $X_{j}$, this additional information would be zero. In general, every $\Delta\mathcal{T}^{i\to j}_{[k,k+1]}$ is a mutual information and hence nonnegative.
Figure 1: (a) Schematic representation of the increase in transfer entropy from variable $X_{i}$ to $X_{j}$ at the $(k+1)$-th timestep. Other variables $X_{l}$ may mediate information transfer even in the absence of a direct coupling from $X_{i}$ to $X_{j}$. (b) Propagation of reference dynamics for $X_{i}$ and $X_{l}$ such that it is commensurate with the given frozen $X_{j}^{(\nu)}$ trajectory. (c) In the RR scheme, trajectories sampled from a reference distribution $P_{0}(X_{i,[0,k]},X_{l,[0,k]})$ are resampled periodically to turn them into the desired conditional distribution $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]}^{(\nu)})$.
For calculating the transfer entropy using Eq. 4 we develop TE-PWS. Following PWS18, the central idea is that trajectory likelihoods in the full $d$-dimensional space are analytically available on-the-fly, and that trajectory averages can be computed in a Monte-Carlo fashion. First, the average in Eq. 4 is computed as,
$\mathcal{T}_{X_{i}\to X_{j}}$ $=\frac{1}{M_{1}}\sum_{\nu}\sum_{k}\ln\frac{P\left(X^{(\nu)}_{j}(k+1)\big{|}X^{(\nu)}_{i,[0,k]},X^{(\nu)}_{j,[0,k]}\right)}{P\left(X^{(\nu)}_{j}(k+1)\big{|}X^{(\nu)}_{j,[0,k]}\right)}$ (5)
where the index $\nu$ sums over $M_{1}$ pairs of trajectories of $X_{i}$ and $X_{j}$ sampled from the joint probability distribution $P\left(X_{i,[0,N]}^{(\nu)},X_{j,[0,N]}^{(\nu)}\right)$. For each pair of trajectories, the probabilties in the numerator and the denominator of Eq. 5 are not analytically available, but what is indeed available is the full joint probability $P(\mathbf{X}_{[0,N]}^{(\nu)})$ as the exponential of the Onsager-Machlup action20, 21. We thus need to marginalize over all degrees of freedom other than $X_{i}$ and $X_{j}$, denoted henceforth collectively as $X_{l}$. We illustrate this procedure first for the denominator in Eq. 5. It is obtained as $P(X^{(\nu)}_{j,[0,k+1]})/P(X^{(\nu)}_{j,[0,k]})$, where $P(X^{(\nu)}_{j,[0,k]})$ is obtained via marginalization,
$P\left(X_{j,[0,k]}^{(\nu)}\right)$
$=\int\int D[X_{i,[0,k]}]D[X_{l,[0,k]}]P\left(X_{i,[0,k]},X_{j,[0,k]}^{(\nu)},X_{l,[0,k]}\right)$ (6)
For performing this average in a Monte-Carlo fashion19, we sample from a reference distribution, $P_{0}(X_{i,[0,k]},X_{l,[0,k]})$, and correct the resultant bias by dividing by $P_{0}$,
$P\left(X_{j,[0,k]}^{(\nu)}\right)$ $=\frac{1}{M_{2}}\sum_{\mu}\frac{P\left(X_{i,[0,k]}^{(\mu)},X_{j,[0,k]}^{(\nu)},X_{l,[0,k]}^{(\mu)}\right)}{P_{0}\left(X_{i,[0,k]}^{(\mu)},X_{l,[0,k]}^{(\mu)}\right)}$ (7)
where the index $\mu$ sums over $M_{2}$ trajectories sampled from $P_{0}$. What is the best choice for $P_{0}$? The ideal choice would be the conditional distribution $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]}^{(\nu)})$, as it makes the summand in Eq. 7 equal for all $\mu$, such that the variance of the estimate of $P(X_{j,[0,k]}^{(\nu)})$ is zero. However, this conditional distribution is not known a priori. We therefore generate $X_{i}$ and $X_{l}$ trajectories in the frozen field of $X_{j,[0,N]}^{(\nu)}$ (Fig. 1b) resulting in a distribution $P_{0}$ that is known analytically and is expected to be close to this conditional distribution. To exactly compensate for the remaining deviations of $P_{0}$ from the desired conditional distribution, we employ, in the spirit of Rosenbluth-Rosenbluth(RR)-PWS18, a reweighing of the $X_{i}^{(\mu)}$ and $X_{l}^{(\mu)}$ trajectories on-the-fly with weights proportional to the ratio of the two distributions. We resample the trajectories with these weights after every $\delta t$ time, meaning that we sample $M_{2}$ trajectories with replacement from the weighted trajectory ensemble (Fig. 1c). This procedure exactly generates $X_{i}^{(\mu)}$ and $X_{l}^{(\mu)}$ trajectories according to the desired conditional distribution $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]}^{(\nu)})$ (see End Matter (EM) and Supplemental Material (SM)).
Returning now to the numerator in Eq. 5, we note that it is an average over another conditional distribution,
$\hskip 14.45377ptP\left(X_{j}^{(\nu)}(k+1)\big{|}X_{i,[0,k]}^{(\nu)},X_{j,[0,k]}^{(\nu)}\right)$
$=\int D[X_{l,[0,k]}]\;P\left(X_{j}^{(\nu)}(k+1)\big{|}X_{i,[0,k]}^{(\nu)},X_{j,[0,k]}^{(\nu)},X_{l,[0,k]}\right)$
.$\qquad\qquad\qquad\;.\;P\left(X_{l,[0,k]}\big{|}X_{i,[0,k]}^{(\nu)},X_{j,[0,k]}^{(\nu)}\right)$ (8)
The first probability in the integral, which is the transition probability of $X_{j}$ in the full $d$-dimensional space, is analytically available. Additionally, similar to the procedure for the denominator in Eq. 5, samples from $P(X_{l,[0,k]}|X_{i,[0,k]}^{(\nu)},X_{j,[0,k]}^{(\nu)})$ are also available by sampling $X_{l}$ trajectories first from a $P_{0}(X_{l,[0,k]})$, and then applying the RR scheme. Thus the numerator in Eq. 5 can also be evaluated in a Monte-Carlo fashion (see EM).
To summarize, for each of the $M_{1}$ pairs of $X_{i,[0,N]}$ and $X_{j,[0,N]}$ trajectories, we simulate a joint ensemble of $M_{2}$ new $X_{i}$ and $X_{l}$ trajectories to estimate the denominator in the logarithm in Eq. 5, and a separate ensemble of $M_{2}$ new $X_{l}$ trajectories to estimate the numerator (Fig. 1). The trajectories in each ensemble are resampled on-the-fly with the RR scheme, giving both the numerator and denominator in Eq. 5 as Monte-Carlo averages. The computational cost thus scales as $2M_{1}M_{2}$. The transfer entropy estimate is unbiased 18 and the statistical accuracy can be arbitrarily improved by increasing $M_{1}$ and $M_{2}$. A pseudocode for the algorithm is available in the SM. Aside from Schreiber’s transfer entropy1, other trajectory-based metrics of directional information transfer such as directed information22, conditional transfer entropy23 and filtered transfer entropy24, also can be derived from conditional distributions of trajectories. Hence TE-PWS can be used to compute all such metrics at similar cost, as shown in the SM.
We demonstrate the validity of the method by analyzing two simple examples of an Ornstein-Uhlenbeck (OU) process for which transfer entropy rates are analytically available. Eq. 1 is an OU process when it is linear with $\mathbf{F}=-\mathbf{a}\mathbf{X}$ where $\mathbf{a}$ is a spring constant matrix. The first example is a minimal three-dimensional model for describing the stochastic dynamics of gene expression and the growth rate of bacteria25. The model is non-bipartite, meaning that the diffusion constant matrix is non-diagonal and the mutual information rates are not finite. Yet transfer entropy rates are finite and semi-analytically available24. The other example is an OU process with two variables, for which transfer entropy rates are analytically available (see EM). As shown in Fig. 2a and b, TE-PWS results converge to the known transfer entropy rates within a few multiples of the relaxation timescales in the respective systems. Details about the models are provided in the EM.
Figure 2: Convergence of transfer entropy rate estimates. (a) and (b) respectively show forward and reverse transfer entropy rates in the stochastic gene expression model in units of an inverse relaxation timescale, $\mu_{0}^{-1}$, as a function of the trajectory duration $T$. Symbols are from TE-PWS and dashed lines are numerically exact values. (c) and (d) show transfer entropy rates in the two-variable OU process. Symbols are from TE-PWS and dashed lines are analytical results.
Jump process. For systems with jumps between a discrete number of states and finite waiting times between the jumps, such as a well-stirred chemical reaction network described by a master equation, or a neural spiking process, the dynamics is governed by a jump propensity matrix $\mathbf{Q}$ of dimensions $\mathcal{N}^{d}\times\mathcal{N}^{d}$, which describes jumps among the $\mathcal{N}$ states of each of the $d$ components of $\mathbf{X}$ 26. In contrast with diffusive processes which must be time-discretized to simulate, jump processes can be simulated exactly with event-driven kinetic Monte-Carlo algorithms such as the Gillespie algorithm27. Hence the transfer of information from $X_{i}$ to $X_{j}$ during the $(k+1)$-th trajectory segment occurs at all instants of time within the segment $X_{j,[k,k+1]}$, rather than at only the endpoint $X_{j}(k+1)$. Therefore in the definition of the stepwise increments to transfer entropy, $\Delta\mathcal{T}_{[k,k+1]}^{i\to j}$, in Eq. 4, the logarithms of the probabilities $P(X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]})$ and $P(X_{j}(k+1)|X_{j,[0,k]})$ should be replaced with functionals of the entire $X_{i,[k,k+1]}$ and $X_{j,[k,k+1]}$ segments,
[…]
$\Delta\mathcal{T}_{[k,k+1]}^{i\to j}$ $=\left\langle\pi_{X_{i}\to X_{j}}-\pi_{X_{j}}\right\rangle$ (11)
where $\alpha$ counts the jumps that change the state of $X_{j}$, $\lambda_{ij}$ and $\lambda_{j}$ are escape propensities for $X_{j}$ in the marginal spaces of $(X_{i},X_{j})$ and $(X_{j})$ respectively, and $\mathcal{Q}_{ij}(\alpha)$ and $\mathcal{Q}_{j}(\alpha)$ similarly are marginal jump propensities for the $\alpha$-th jump2, 15. $Q_{ij},\lambda_{ij}$ and $Q_{j},\lambda_{j}$ are obtained by marginalizing jump propensities from the full $\mathbf{X}$-space into the $(X_{i},X_{j})$ and $(X_{j})$ spaces respectively2, 15 (see SM).
The transfer entropy in Eq. 11 can be calculated with TE-PWS by marginalizing jump propensities in a Monte-Carlo fashion over conditional distributions of hidden variables, similar to the diffusive case. Specifically, the average in Eq. 11 is expressed as a Monte-Carlo average over trajectories of $X_{i}$ and $X_{j}$, and each of $\mathcal{Q}_{ij},\lambda_{ij}$ and $\mathcal{Q}_{j},\lambda_{j}$ is expressed as a Monte-Carlo average over conditional distributions of hidden variables. Hidden variable trajectories are sampled from a $P_{0}$ propagated in the field of the frozen trajectory, and the remaining bias is exactly corrected for through the RR scheme (see SM). Prior work argues that the escape terms involving $\lambda_{ij}$ and $\lambda_{j}$ in Eqs. 9 and 10 respectively can be omitted since they cancel each other on average2. We show however in the SM that the error in the transfer entropy estimate can be reduced by an order of magnitude by exploiting anti-correlated fluctuations between the escape and the jump terms.
Data processing inequality. We demonstrate the utility of TE-PWS by applying it to a three-node motif to study whether transfer entropies obey the Data Processing Inequality (DPI). Unidirectional flow of information between different nodes in a network leads to a DPI for the mutual information28. For a general three-variable process, if the flow of information is $X_{1}\to X_{2}\to X_{3}$, i.e., without feedback, $I(X_{1,[0,N]};X_{3,[0,N]}|X_{2,[0,N]})=0$; here the right arrows denote flow of information as mediated either via activation or repression. This leads by the chain rule to $I(X_{1,[0,N]};X_{3,[0,N]})\leq I(X_{1,[0,N]};X_{2,[0,N]})$28. As transfer entropy equals mutual information in the absence of feedback, it also obeys $\mathcal{T}_{X_{1}\to X_{3}}\leq\mathcal{T}_{X_{1}\to X_{2}}$. This relation bounds the amount of information that can be transmitted from input to output through an intermediate variable, yet is only valid in the absence of feedback. In the presence of an $X_{2}\rightarrow X_{1}$ feedback, i.e., $X_{1}\rightleftarrows X_{2}\to X_{3}$, the mutual information continues to obey its DPI, while the transfer entropy formally doesn’t29, 30. This can be rationalized by considering the limit $X_{1}\leftarrow X_{2}\rightarrow X_{3}$, where $X_{2}$ controls both $X_{1}$ and $X_{3}$. Here the $X_{1}$ trajectories would still be predictive of fluctuations in $X_{3}$, thus $\mathcal{T}_{X_{1}\to X_{3}}\geq 0$ even as $\mathcal{T}_{X_{1}\to X_{2}}=0$. However, where the crossover from the feedforward to the feedback-dominated regime occurs, and to what extent transfer entropies can overcome the DPI in practice, is currently not understood.
Figure 3: (a) and (b) are typical $X_{1}$ (green) and $X_{2}$ (black) trajectories for models A and B at $a_{12}=2a_{21}=0.4$ and $a_{12}=a_{21}=-6$ respectively. (c) Transfer entropy rates in model A, denoted as $\dot{\mathcal{T}}_{X_{1}\to X_{3}}^{(\mathrm{A})}$ and $\dot{\mathcal{T}}_{X_{1}\to X_{2}}^{(\mathrm{A})}$, as a function of an increasing ratio of feedback to feedforward strength $f^{*}=a_{12}/a_{21}$, where the feedforward strength is kept constant at $a_{21}=0.2$. (d) Similar to (c) but for model B, where the feedforward strength is kept constant at $a_{21}=-4$. (e) Ratio of transfer entropy rates overcomes DPI bound (dashed line at $\mathcal{T}^{*}=1$) with increasing feedback.
We implement the motif $X_{1}\rightleftarrows X_{2}\to X_{3}$ in two diffusive models of mutual repression between $X_{1}$ and $X_{2}$, labeled as models A and B. In both models, $X_{3}$ rapidly copies the state of $X_{2}$ such that the information loss from $X_{2}$ to $X_{3}$ is low. Model A is a three-dimensional OU process with linear feedback, where the ratio of the feedback to feedforward spring constants $f^{*}\equiv a_{12}/a_{21}$ is varied to study the violation of DPI. Model B is a nonlinear extension of A inspired by a genetic toggle switch31 where the drifts for $X_{1}$ and $X_{2}$ are changed to $F_{1}=-a_{11}X_{1}-a_{12}(1+X_{1}^{2})/(1+X_{1}^{2}+X_{2}^{2})$ and $F_{2}=-a_{22}X_{2}-a_{21}(1+X_{2}^{2})/(1+X_{1}^{2}+X_{2}^{2})$; we choose $a_{12}/a_{21}=f^{*}$ such that the two models can be compared. Other model parameters are provided in the EM. Typical $X_{1}$ and $X_{2}$ trajectories at high feedback are shown in Figs. 3a and b. Model A only shows regression to the mean for both $X_{1}$ and $X_{2}$, while model B additionally shows switching between (low,high) and (high,low) values of $(X_{1},X_{2})$ in the regime of both strong feedforward and feedback coupling, around $f^{*}=1$.
We plot in Figs. 3c and d transfer entropy rates for the two models. Each data point is obtained from steady state trajectories using TE-PWS. In model A, increasing the feedback strength $f^{*}$ keeps $\dot{\mathcal{T}}_{X_{1}\to X_{2}}$ unchanged while $\dot{\mathcal{T}}_{X_{1}\to X_{3}}$ monotonically increases. The former follows analytically from the expression for the transfer entropy in a two-dimensional OU process24 (see EM). The latter is an empirical result that arises because, when $f^{*}$ is increased, $X_{1}$ becomes increasingly correlated with $X_{2}$ while $X_{2}$ continues to be copied accurately by $X_{3}$. In contrast, in model B, both $\dot{\mathcal{T}}_{X_{1}\to X_{2}}$ and $\dot{\mathcal{T}}_{X_{1}\to X_{3}}$ peak near the switching regime, phenomenologically similar to how a bistable three-node motif with feedback has been shown to behave through approximate theory15. The values of the transfer entropy rates in model B are also amplified many-fold compared to model A due to stronger correlations between $X_{1}$ and $X_{2}$ resulting from the stronger, nonlinear coupling, with the concomitant switching.
In Fig. 3e we have plotted the amount of DPI violation in both models, quantified as the ratio of transfer entropy rates $\mathcal{T}^{*}\equiv\dot{\mathcal{T}}_{X_{1}\to X_{3}}/\dot{\mathcal{T}}_{X_{1}\to X_{2}}$, as a function of the ratio of the feedback to feedforward strengths, $f^{*}$. Surprisingly, we find that regardless of the nature of the variation of the individual transfer entropy rates with increasing feedback, i.e., monotonic or non-monotonic (Figs. 3c and d), the ratio $\mathcal{T}^{*}$ monotonically increases with increasing feedback (Fig. 3 e). Moreover, the ratio overcomes the DPI bound when the strength of the feedback becomes comparable to that of the feedforward coupling. Our results thus show that when the feedback $X_{2}\to X_{1}$ dominates over the feedforward interaction, $X_{1}\to X_{2}$, the feedforward entropy $\mathcal{T}_{X_{1}\to X_{3}}$ becomes larger than the feedforward entropy $\mathcal{T}_{X_{1}\to X_{2}}$. The mutual information, on the other hand, continues to obey its DPI, $I(X_{1,[0,N]};X_{3,[0,N]})\leq I(X_{1,[0,N]};X_{2,[0,N]})$. We expect further analytical work in the OU process to be able to support this empirical result.
In conclusion, we have developed a method that for the first time makes it possible to compute transfer entropies exactly for any stochastic model. We have shown that TE-PWS is applicable to diverse stochastic processes and can elucidate novel physics of information transmission. We expect transfer entropies computed by TE-PWS to be used as ground truth for a wide range of goals, such as the characterization and design of information flow in natural and engineered information processing systems, and causality detection.
Acknowledgement We thank Manuel Reinhardt and Age Tjalma for useful discussions and Vahe Galstyan for a careful reading of the manuscript. This work is part of the Dutch Research Council (NWO) and was performed at the research institute AMOLF. This project has received funding from the European Research Council under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 885065).
Data availability Python implementations of TE-PWS for diffusive and jump processes, and data that reproduce the findings of this study, are openly available on Zenodo at https://zenodo.org/doi/10.5281/zenodo.1361736532.
Appendix A End Matter
Model details. Figs. 2 a and b show transfer entropies in a minimal stochastic model for gene expression25, 11, 24. The model is a three-dimensional OU process with spring constants $a_{11}=\mu_{E}+\mu_{0}T_{\mu E}(T_{EG}-1)$, $a_{12}=-\mu_{0}(T_{EG}-1)$, $a_{13}=\mu_{0}T_{EG}$, $a_{21}=-T_{\mu E}[\beta_{G}-\mu_{E}-\mu_{0}T_{\mu E}(T_{EG}-1)]$, $a_{22}=-[\mu_{0}T_{\mu E}(T_{EG}-1)-\beta_{G}]$, $a_{23}=-[\beta_{G}-\beta_{\mu}-\mu_{0}T_{\mu E}T_{EG}]$, $a_{31}=a_{32}=0$ and $a_{33}=\beta_{\mu}$, and diffusion constants $D_{11}=D_{E}\mu_{0}^{2}$, $D_{22}=\beta_{\mu}\eta_{\mu}^{2}+\beta_{G}\eta_{G}^{2}+D_{E}\mu_{0}^{2}T_{\mu E}^{2}$, $D_{33}=\beta_{\mu}\eta_{\mu}^{2}$, $D_{12}=D_{21}=D_{E}\mu_{0}^{2}T_{\mu E}$, $D_{13}=D_{31}=0$ and $D_{23}=D_{32}=\beta_{\mu}\eta_{\mu}^{2}$, where the experimentally determined values of the parameters are $\mu_{0}=0.23h^{-1}$, $\beta_{\mu}=\beta_{G}=0.33h^{-1}$, $\beta_{E}=5.63h^{-}1$, $\eta_{E}=1.03$, $\eta_{\mu}=0.16$, $\eta_{G}=0.22$, $T_{EE}=1$, $T_{\mu E}=0.7$, $T_{EG}=1.3$, $\mu_{E}=\mu_{0}(1+T_{\mu E}-T_{EE})$ and $D_{E}=\eta_{E}^{2}/\beta_{E}$.25, 24. We also used $\delta t=9.2\times 10^{-5}\mu_{0}^{-1}$, $M_{1}=96$ and $M_{2}=200$.
For Figs. 2c and d, the model is a two-dimensional OU process with parameters $a_{11}=1$, $a_{22}=2$, $a_{12}=a_{21}=0.5$, $D_{11}=1.5$, $D_{22}=0.5$ and $D_{12}=0$. Additionally, $\delta t=0.01a_{11}^{-1}$, $M_{1}=100$ and $M_{2}=1000$.
For models A and B in Fig. 3, we have chosen $a_{11}=a_{22}=1$, $a_{33}=-a_{32}=2$, $a_{13}=a_{31}=a_{23}=0$ and $D_{ij}=\delta_{ij}$. Additionally, $\delta t=0.01a_{11}^{-1}$, $M_{1}=100$ and $M_{2}=1000$.
In all simulations in the main text, trajectories have been initialized from a histogram of the corresponding steady-state distributions. Error bars for all results are computed as twice the standard deviations using multiple sets of statistically independent simulations.
Analytical expression in two-dimensional OU process. In a two-dimensional OU process $(X_{1},X_{2})$ with spring constants $a_{ij}$ and diffusion constant constants $D_{ij}$ for $i,j\in\{1,2\}$, the steady-state transfer entropy rate $\dot{\mathcal{T}}_{X_{1}\to X_{2}}$ is given by24
$\dot{\mathcal{T}}_{X_{1}\to X_{2}}=\frac{1}{2}\left(r_{2}-a_{11}+\frac{D_{12}}{D_{22}}a_{21}\right)$ (12)
where $r_{2}=[a_{11}^{2}+(D_{11}/D_{22})a_{21}^{2}-2(D_{12}/D_{22})a_{11}a_{21}]^{1/2}$. For model A in the section Data processing inequality of the main text, the dynamics of $X_{1}$ and $X_{2}$ comprises that of a two-dimensional OU process because there is no feedback from $X_{3}$. Hence the transfer entropy rate $\dot{\mathcal{T}}_{X_{1}\to X_{2}}$ is given by Eq. 12, which demonstrably does not depend on the $X_{2}\to X_{1}$ feedback strength $a_{12}$, i.e., on $f^{*}$ (see Fig. 3c).
TE-PWS algorithm. Central to TE-PWS is the computation of trajectory averages in a Monte-Carlo fashion over simulated trajectories, and the availability of trajectory probabilities on-the-fly in the full $d$-dimensional space. To implement TE-PWS for a given stochastic model, we need to specify the method to simulate the model and an explicit functional form of the trajectory probability.
The diffusion processes modeled by Eq. 1 can be simulated using an Euler-Maruyama scheme with a fixed small timestep. In our examples we have taken the timestep to be equal to the duration of the trajectory segments for implementing the RR scheme, $\delta t$, for convenience. The propagation equation for the $(k+1)$-th step is $\mathbf{X}(k+1)=\mathbf{X}(k)+\delta t\mathbf{F}(k)+\sqrt{\delta t}\bm{\psi}(k)$, where $\bm{\psi}$ is a Gaussian random vector with zero mean and variance $\langle\psi_{i}(k)\psi_{j}(k^{{}^{\prime}})\rangle=2D_{ij}\delta_{k,k^{{}^{\prime}}}$. Making time discrete results in an $\mathcal{O}(\delta t)$ error which can be made arbitrarily small by systematically decreasing the timestep. The probability density of the change of state $\Delta\mathbf{X}(k)$ can be written analytically through Ito discretization of the Onsager-Machlup action as 20, 21
$P(\Delta\mathbf{X}(k))=$ $\frac{1}{(4\pi\delta t)^{d/2}|\mathbf{D}|}\exp\big{[}-(\Delta\mathbf{X}(k)-\mathbf{F}(k)\delta t)^{T}$
$\qquad\cdot\mathbf{D}^{-1}(\Delta\mathbf{X}(k)-\mathbf{F}(k)\delta t)/4\delta t\big{]}$ (13)
where $|\mathbf{D}|$ is the determinant of the diffusion constant matrix. This form also holds for systems with inertia if the generalized coordinate vector $\mathbf{X}$ contains both positions and velocities33, 34.
Jump processes can be simulated with a Gillespie algorithm, which is exact, i.e., does not make a timestep error27. To clarify the quantities in the main text (Eqs. 9-11), and elucidate how they can be obtained in the Gillespie simulations, we briefly describe the scheme. At any given time, the waiting time till the next jump is exponentially distributed. Hence at the start of the trajectory and after every jump, denoting the time as $t$, we sample a uniform random number $u\in[0,1]$ and determine the wait time $\tau$ by solving
[…]
for $\tau$. Here $\lambda(t^{{}^{\prime}})$ is the escape propensity from state $\mathbf{X}(t^{{}^{\prime}})$, given by the sum of the jump propensities $\mathcal{Q}_{\beta}(t^{{}^{\prime}})$ taking the system out of the state $\mathbf{X}(t^{{}^{\prime}})$ at time $t^{{}^{\prime}}$. When propagating all coordinates at once, $\lambda(t^{{}^{\prime}})$ is independent of $t^{{}^{\prime}}$ in between jumps; hence Eq. 14 is trivially solved for $\tau$. Then a second uniform random number is drawn to choose which jump to fire, with probabilities proportional to their jump propensities. The probability density of a trajectory segment $\mathbf{X}_{[t,t+\delta t]}$ is written as
[…]
where $\alpha$ sums over all jumps in the full trajectory, $N_{\mathrm{tot}}$ in number, occurring at times $t_{\alpha}$.
The elementary steps of the TE-PWS algorithm are the same for both diffusive and jump processes. In brief:
  1. 1.
    Propagate $M_{1}$ trajectories of $X_{i}$ and $X_{j}$ in the full $d$-dimensional space. These trajectories are henceforth labeled with $(\nu)$.
    • Steps 2-5 are for computing the numerator in Eq. 5 for diffusive processes and the first term in Eq. 11 for jump processes.
  2. 2.
    For each pair of $(X_{i}^{(\nu)},X_{j}^{(\nu)})$ trajectories, propagate $M_{2}$ trajectories of hidden variables $X_{l}$ using the chosen reference distribution $P_{0}(X_{l,[0,k]})$. Initial conditions should be sampled from the same joint distribution $P(X_{i}^{(\nu)}(0),X_{j}^{(\nu)}(0),X_{l}(0))$ that $X_{i}^{(\nu)}(0)$ and $X_{j}^{(\nu)}(0)$ were drawn from.
  3. 3.
    After every $\delta t$ time, recalculate logarithmic weights $w^{(\mu)}$ for the trajectories defined as the logarithm of the ratio between the joint distribution $P(X_{i,[0,k]}^{(\nu)},X_{j,[0,k]}^{(\nu)},X_{l,[0,k]})$ and the reference distribution $P_{0}(X_{l,[0,k]})$, using Eqs. A and 14. This ratio is proportional to the ratio between the conditional distribution $P(X_{l,[0,k]}|X_{i,[0,k]}^{(\nu)},X_{j,[0,k]}^{(\nu)})$ and the reference distribution $P_{0}(X_{l,[0,k]})$, as proven in the SM. Then calculate the uniformity in the weights with a uniformity parameter $\kappa=(\sum_{\mu}\exp w^{(\mu)})^{2}/\sum_{\mu}\exp(2w^{(\mu)})$, where sums of exponentials of weights are always performed with the Log-Sum-Exp trick35.
  4. 4.
    Calculate the contribution to the transfer entropy in the space of $(X_{i},X_{j})$ at the $(k+1)$-th step, denoted as $\mathcal{T}_{a}^{(\nu)}[k]$. This is done by computing the average in Eq. 8 for diffusive processes, and taking an expectation of Eq. 9 for jump processes (see SM).
  5. 5.
    If $\kappa<M_{2}/2$, resample the $M_{2}$ trajectories with the accumulated weights $w^{(\mu)}$ (see Fig. 1c). This means we sample $M_{2}$ trajectories randomly with weights $w^{(\mu)}$ from the simulated trajectory ensemble, with replacement. Set all weights to zero after every resampling.
    • Step 6 is for computing the denominator in Eq. 5 for diffusive processes and the second term in Eq. 11 for jump processes.
  6. 6.
    Akin to steps 2-5, for each $(X_{j}^{(\nu)})$ trajectory, propagate $M_{2}$ trajectories of $(X_{i},X_{l})$ using reference distribution $P_{0}(X_{i,[0,k]},X_{l,[0,k]})$; accumulate weights for each trajectory as the ratio between the joint distribution $P(X_{i,[0,k]},X_{j,[0,k]}^{(\nu)},X_{l,[0,k]})$ and the reference distribution $P_{0}(X_{i,[0,k]},X_{l,[0,k]})$; compute the contribution to the transfer entropy in the $(X_{j})$ space at the $(k+1)$-th step, denoted as $\mathcal{T}_{b}^{(\nu)}[k]$ (Eqs. 7 and 10); and resample if the uniformity in the weights is low, i.e., $\kappa<M_{2}/2$.
  7. 7.
    Finally, compute $\mathcal{T}_{X_{i}\to X_{j}}$ by combining $\mathcal{T}_{a}^{(\nu)}$ and $\mathcal{T}_{b}^{(\nu)}$ from all timesteps using Eqs. 5 and 11 for diffusive and jump processes respectively.
For clarity, a pseudocode to compute transfer entropy $\mathcal{T}_{X_{1}\to X_{3}}$ in a three-variable process is given in the SM. We note that TE-PWS is easily parallelized as it is a Monte-Carlo algorithm. Specifically, steps 1-7 can be executed independently for each of the $M_{1}$ trajectories mentioned in step 1, using parallel processors. The transfer entropy is computed at the end as an average over the results from all processors.
Supplemental Material:
Exact computation of Transfer Entropy with Path Weight Sampling
November 25, 2024
Appendix B Pseudocode
A pseudocode for computing the transfer entropy $\mathcal{T}_{X_{1}\to X_{3}}$ in a three-variable stochastic process via the TE-PWS algorithm is provided in Algorithm 1.
Algorithm 1 Transfer Entropy-Path Weight Sampling (TE-PWS)
1:objective computation of the transfer entropy $\mathcal{T}_{X_{1}\to X_{3}}$ in a three-variable stochastic process.
2:inputs model parameters for simulating $\mathbf{X}(t)$
3:parameters timestep $\delta t$; number of timesteps $N$; number of trajectories for Monte-Carlo averages $M_{1}$ and $M_{2}$; resampling indicator $\kappa$
4:initialize Define trajectory labels $\nu$ and $\mu$; timestep variable $k$; cumulative transfer entropy array $\mathcal{T}[0:N]$; two terms (factors) of transfer entropy $\mathcal{T}_{a}^{(\nu)}[0:N]$ and $\mathcal{T}_{b}^{(\nu)}[0:N]$ respectively.
5:$\nu\leftarrow 0$
[…]
7:repeat
8:     Generate $M_{1}$ trajectories $(X_{1,[0,N]}^{(\nu)},X_{2,[0,N]}^{(\nu)},X_{3,[0,N]}^{(\nu)})$ jointly.
9:     $k\leftarrow 0$triangleright Accessing $P(X_{2,[0,N]}|X_{1,[0,N]},X_{3,[0,N]})$
10:     $\kappa\leftarrow M_{2}$
11:     initialize Generate $M_{2}$ samples of initial conditions $X_{2}^{(\mu)}(0)$ labeled by $\mu$ from a steady-state trajectory; weights in log scale $w^{(\mu)}=0$.
12:     repeat
13:         if $\kappa<M_{2}/2$ then
14:              Resample $M_{2}$ configurations from $X_{2}^{(\mu)}(k)$ with weights $\exp w^{(\mu)}$.
15:              $w^{(\mu)}\leftarrow 0$ for all $\mu$
16:         end if
17:         Propagate reference dynamics $X_{2,[k,k+1]}^{(\mu)}$ as samples from $P_{0}(X_{2,[k,k+1]}|X_{2,[0,k]})$.
18:         $w^{(\mu)}\leftarrow w^{(\mu)}+\ln P\left(X_{1,[k,k+1]}^{(\nu)},X_{2,[k,k+1]}^{(\mu)},X_{3,[k,k+1]}^{(\nu)}\big{|}X_{1,[0,k]}^{(\nu)},X_{2,[0,k]}^{(\mu)},X_{3,[0,k]}^{(\nu)}\right)-\ln P_{0}\left(X_{2,[k,k+1]}^{(\mu)}\big{|}X_{2,[0,k]}^{(\mu)}\right)$
19:         $\kappa\leftarrow(\sum_{\mu}\exp w^{(\mu)})^{2}/\sum_{\mu}\exp(2w^{(\mu)})$
20:         Compute $\mathcal{T}_{a}^{(\nu)}[k]$ using Eq. 8 or 9 of the main text for a diffusion or jump process respectively.
21:         $k\leftarrow k+1$
22:     until $k=N$
23:     $k\leftarrow 0$triangleright Accessing $P(X_{1,[0,N]},X_{2,[0,N]}|X_{3,[0,N]})$
24:     $\kappa\leftarrow M_{2}$
25:     initialize Generate $M_{2}$ samples of initial conditions $(X_{1}^{(\mu)}(0),X_{2}^{(\mu)}(0))$ labeled by $\mu$ from a steady-state trajectory; weights in log scale $w^{(\mu)}=0$.
26:     repeat
27:         if $\kappa<M_{2}/2$ then
28:              Resample $M_{2}$ configurations from $(X_{1}^{(\mu)}(k),X_{2}^{(\mu)}(k))$ with weights $\exp w^{(\mu)}$.
29:              $w^{(\mu)}\leftarrow 0$ for all $\mu$
30:         end if
31:         Propagate reference dynamics $(X_{1,[k,k+1]}^{(\mu)},X_{2,[k,k+1]}^{(\mu)})$ as samples from $P_{0}(X_{1,[k,k+1]},X_{2,[k,k+1]}|X_{1,[0,k]},X_{2,[0,k]})$.
32:         $w^{(\mu)}\leftarrow w^{(\mu)}+\ln P\left(X_{1,[k,k+1]}^{(\mu)},X_{2,[k,k+1]}^{(\mu)},X_{3,[k,k+1]}^{(\nu)}\big{|}X_{1,[0,k]}^{(\mu)},X_{2,[0,k]}^{(\mu)},X_{3,[0,k]}^{(\nu)}\right)$
33:         $w^{(\mu)}\leftarrow w^{(\mu)}-\ln P_{0}\left(X_{1,[k,k+1]}^{(\mu)},X_{2,[k,k+1]}^{(\mu)}\big{|}X_{1,[0,k]}^{(\mu)},X_{2,[0,k]}^{(\mu)}\right)$
34:         $\kappa\leftarrow(\sum_{\mu}\exp w^{(\mu)})^{2}/\sum_{\mu}\exp(2w^{(\mu)})$
35:         Compute $\mathcal{T}_{b}^{(\nu)}[k]$ using Eq. 7 or 10 in the main text for a diffusion or jump process respectively.
36:         $k\leftarrow k+1$
37:     until $k=N$
38:     $\mathcal{T}[0:N]\leftarrow\mathcal{T}[0:N]+\mathcal{T}^{(\nu)}_{a}[0:N]-\mathcal{T}^{(\nu)}_{b}[0:N]$
39:     $\nu\leftarrow\nu+1$
40:until $\nu=M_{1}$
[…]
Appendix C Proof of the efficiency of the RR scheme
Here we prove that the Rosenbluth-Rosenbluth (RR) scheme in TE-PWS achieves optimal efficiency for a Monte-Carlo evaluation of the transfer entropy. For clarity, we start by reproducing Eq. 7 of the main text here, representing a typical marginalization procedure in TE-PWS,
$P\left(X_{j,[0,k]}^{(\nu)}\right)$ $=\frac{1}{M_{2}}\sum_{\mu}\frac{P\left(X_{i,[0,k]}^{(\mu)},X_{j,[0,k]}^{(\nu)},X_{l,[0,k]}^{(\mu)}\right)}{P_{0}\left(X_{i,[0,k]}^{(\mu)},X_{l,[0,k]}^{(\mu)}\right)}$ (S1)
TE-PWS performs this average with high statistical efficiency by preferentially sampling rare large values of the summand. Similar to PWS as discussed in [1], this is implemented using the RR scheme, in which the ensemble of $M_{2}$ trajectories is resampled after every $\delta t$ time (see Fig. 1 of main text). For example, in order to compute $P(X_{j,[0,N]}^{(\nu)})$, an ensemble of $X_{i}$ and $X_{l}$ trajectories is simulated. At the $(k+1)$-th step, the resampling weight used for the $\mu$-th trajectory in the ensemble is
$\widehat{g}\big{[}X_{l,[0,k+1]}^{(\mu)}\big{]}=P\left(X_{i,[k,k+1]}^{(\mu)},X_{j,[k,k+1]}^{(\nu)},X_{l,[k,k+1]}^{(\mu)}\Big{|}X_{i,[0,k]}^{(\mu)},X_{j,[0,k]}^{(\nu)},X_{l,[0,k]}^{(\mu)}\right)/\;P_{0}\left(X_{i,[k,k+1]}^{(\mu)},X_{l,[k,k+1]}^{(\mu)}\Big{|}X_{i,[0,k]}^{(\mu)},X_{l,[0,k]}^{(\mu)}\right)$ (S2)
which is analytically available. We show below that iteratively resampling the newly generated trajectories with this weight changes the trajectory distribution optimally such that the summand in Eq. S1 is a constant, i.e., has a variance of zero, achieving perfect sampling. We will show this by first showing that if after the $k$-th step the trajectories have an optimal distribution, they stay optimal after the $(k+1)$-th step. Then, combined with the fact that the distribution of initial conditions is by construction optimal, we will conclude by induction that the distribution stays uniform during the entire duration of the trajectory.
The optimal choice for $P_{0}$ would be the hypothetical $P_{0}(X_{i,[0,N]},X_{l,[0,N]})=P(X_{i,[0,N]},X_{l,[0,N]}|X_{j,[0,N]}^{(\nu)})$ as it makes the summand in Eq. S1 a constant independent of the index $\mu$, resulting in a zero-variance estimate of $P(X_{j,[0,N]}^{(\nu)})$. The purpose of the RR scheme is to bias the simulated distribution $P_{0}(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)})$ towards the optimal distribution $P(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)}|X_{j,[0,N]}^{(\nu)})$. For the inductive argument, assume that after the $k$-th step, the trajectories are distributed optimally according to $P(X_{i,[0,k]}^{(\mu)},X_{l,[0,k]}^{(\mu)}|X_{j,[0,k]}^{(\nu)})$. After the next propagation step, the probability of each trajectory changes to a product of its previous value and the probability of the new segment,
$w^{(\mu)}$ $=P\left(X_{i,[0,k]}^{(\mu)},X_{l,[0,k]}^{(\mu)}\Big{|}X_{j,[0,k]}^{(\nu)}\right).\;P_{0}\left(X_{i,[k,k+1]}^{(\mu)},X_{l,[k,k+1]}^{(\mu)}\Big{|}X_{i,[0,k]}^{(\mu)},X_{l,[0,k]}^{(\mu)}\right)$
$=P\left(X_{i,[0,k]}^{(\mu)},X_{j,[0,k]}^{(\nu)},X_{l,[0,k]}^{(\mu)}\right).\;P_{0}\left(X_{i,[k,k+1]}^{(\mu)},X_{l,[k,k+1]}^{(\mu)}\Big{|}X_{i,[0,k]}^{(\mu)},X_{l,[0,k]}^{(\mu)}\right)\big{/}\;P\left(X_{j,[0,k]}^{(\nu)}\right)$ (S3)
Then we resample the trajectories with weight $\widehat{g}$ given in Eq. S2. The probability distribution of these trajectories after this resampling step becomes proportional to the product of Eqs. S2 and S3,
$w^{(\mu)}\widehat{g}\big{[}X_{l,[0,k+1]}^{(\mu)}\big{]}$ $=\frac{P\left(X_{i,[0,k+1]}^{(\mu)},X_{j,[0,k+1]}^{(\nu)},X_{l,[0,k+1]}^{(\mu)}\right)}{P\left(X_{j,[0,k]}^{(\nu)}\right)}$ (S4)
where Bayes’ theorem has been used to condense the numerator. The normalization constant for this probability is obtained by summing Eq. S4 over all $X_{i,[0,k+1]}^{(\mu)}$ and $X_{l,[0,k+1]}^{(\mu)}$ trajectories, which gives $P\left(X_{j,[0,k+1]}^{(\nu)}\right)\big{/}P\left(X_{j,[0,k]}^{(\nu)}\right)$. Dividing Eq. S4 by this normalization constant gives the new normalized probability distribution as $P(X_{i,[0,k+1]}^{(\mu)},X_{l,[0,k+1]}^{(\mu)}|X_{j,[0,k+1]}^{(\nu)})$. Hence the trajectories remain distributed optimally after resampling.
To complete the proof by induction, we also need to show that the initial conditions of the $X_{i,[0,k]}^{(\mu)}$ and $X_{l,[0,k]}^{(\mu)}$ trajectories are consistent with the optimal distribution $P(X_{i,[0,k]}^{(\mu)},X_{l,[0,k]}^{(\mu)}|X_{j,[0,k]}^{(\nu)})$. This requirement is met by construction, because the initial conditions for the $X_{i}^{(\mu)}$ and $X_{l}^{(\mu)}$ trajectories are sampled from the same joint distribution $P(\mathbf{X}(0))$ as that from which $X_{j}^{(\nu)}(0)$ is drawn; if one is interested in the steady-state transfer entropy rates, this distribution is the steady-state joint distribution where $X_{j}=X_{j}^{(\nu)}(0)$ is given. This ingredient, together with the fact that the resampling procedure preserves the optimal distribution as described above, guarantees that, by induction, the RR scheme generates samples from the optimal conditional distribution $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]}^{(\nu)})$ at every step. Resampling is algorithmically performed with a stratified resampling technique which is computationally efficient2. Samples from the optimal distribution are then used in Eq. S1 to compute the denominator in Eq. 5 of the main text, reproduced here,
$\mathcal{T}_{X_{i}\to X_{j}}$ $=\frac{1}{M_{1}}\sum_{\nu}\sum_{k}\ln\frac{P\left(X^{(\nu)}_{j}(k+1)\big{|}X^{(\nu)}_{i,[0,k]},X^{(\nu)}_{j,[0,k]}\right)}{P\left(X^{(\nu)}_{j}(k+1)\big{|}X^{(\nu)}_{j,[0,k]}\right)}$ (S5)
The numerator is obtained similarly by generating samples from the corresponding conditional distribution $P(X_{l,[0,k]}|X_{i,[0,k]}^{(\nu)},X_{j,[0,k]}^{(\nu)})$ with the RR scheme. Conditional distributions of trajectories of all combinations of variables can be similarly sampled by repeating this procedure.
Appendix D Choice of reference distribution
The transfer entropy estimated from TE-PWS is exact for any choice of the reference probability $P_{0}$ because of the RR scheme. However, the number of times resampling needs to be performed depends on how large the variance of the summand in Eq. S1 is. A better choice of $P_{0}$ results in a smaller variance at the same computational cost. As discussed earlier, for computing $P(X_{j,[0,N]}^{(\nu)})$ for example, the ideal choice for $P_{0}$ would be $P_{0}(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)})=P(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)}|X_{j,[0,N]}^{(\nu)})$, which is not known a priori and is impossible to directly sample from. We therefore choose a distribution $P_{0}(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)})$ that uses the past trajectory of $X_{j,[0,N]}^{(\nu)}$ at every timestep to compute the drift and diffusion terms for $X_{i}$ and $X_{l}$, similar to the original dynamics of the system in the full $d$-dimensional space. This keeps the reference distribution close to the target conditional distribution while being analytically known and easy to sample from.
As an example, consider the three-dimensional OU process discussed before, $\dot{\mathbf{X}}=-\mathbf{a}\mathbf{X}+\bm{\xi}$ where $\mathbf{a}$ is the spring-constant matrix and $\bm{\xi}$ is a Gaussian white noise with zero mean and covariance matrix $2\mathbf{D}$. The diffusion constant matrix $\mathbf{D}$ may have nonzero off-diagonal elements. In the full three-dimensional space, propagating the natural dynamics involves computing the drifts and sampling three correlated noise components from the joint Gaussian distribution, which we will call $\mathbb{G}(\xi_{i},\xi_{j},\xi_{l})$. For sampling trajectories $X_{i,[0,N]}^{(\mu)}$ and $X_{l,[0,N]}^{(\mu)}$, for a given $X_{j,[0,N]}^{(\nu)}$, from a distribution $P_{0}(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)})$ that is as close as possible to $P(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)}|X_{j,[0,N]}^{(\nu)})$, we propagate $X_{i}^{(\mu)}$ and $X_{l}^{(\mu)}$ trajectories with the equations of motion
$\dot{X}_{i}^{(\mu)}(k)$ $=-a_{11}X_{i}^{(\mu)}(k)-a_{12}X^{(\nu)}_{j}(k)-a_{13}X_{l}^{(\mu)}(k)+\tilde{\xi}_{i}^{(\mu)}(k)$ (S6)
$\dot{X}_{l}^{(\mu)}(k)$ $=-a_{31}X_{i}^{(\mu)}(k)-a_{32}X^{(\nu)}_{j}(k)-a_{33}X_{l}^{(\mu)}(k)+\tilde{\xi}_{l}^{(\mu)}(k)$ (S7)
where $\tilde{\xi}_{i}^{(\mu)}$ and $\tilde{\xi}_{l}^{(\mu)}$ are Gaussian white noises whose distributions should be commensurate with the existing noises $\xi_{j}^{(\nu)}$ in the $X_{j,[0,N]}^{(\nu)}$ trajectory. However, given only the $X_{j,[0,N]}^{(\nu)}$ trajectory, the noises $\xi_{j}^{(\nu)}(k)$ are not uniquely known since the drifts $F_{j}^{(\nu)}(k)$ are unknown; hence, the distributions of $\tilde{\xi}_{i}^{(\mu)}$ and $\tilde{\xi}_{l}^{(\mu)}$ are not yet uniquely defined. To fully specify these distributions, we now make a choice for deriving approximate noises in $X_{j,[0,N]}^{(\nu)}$, called $\tilde{\xi}_{j}^{(\nu\mu)}$, by assuming that the drift in $X_{j}$ depends, besides $X_{j}^{(\nu)}(k)$ itself, on $X_{i}^{(\mu)}(k)$ and $X_{l}^{(\mu)}(k)$, i.e. the drift is taken to be $-a_{21}X_{i}^{(\mu)}(k)-a_{22}X^{(\nu)}_{j}(k)-a_{23}X_{l}^{(\mu)}(k)$. Next, obtaining $\dot{X}_{j}^{(\nu)}(k)$ from the given $X_{j}^{(\nu)}$ trajectory and with the drift now specified, we solve, at every timestep,
$\dot{X}_{j}^{(\nu)}(k)$ $=-a_{21}X_{i}^{(\mu)}(k)-a_{22}X^{(\nu)}_{j}(k)-a_{23}X_{l}^{(\mu)}(k)+\tilde{\xi}_{j}^{(\nu\mu)}(k)$ (S8)
for $\tilde{\xi}_{j}^{(\nu\mu)}(k)$, which can be done by simply transposing the equation even if the original dynamics was not linear. Now given $\tilde{\xi}_{j}^{(\nu\mu)}(k)$, from which distribution should we generate $\tilde{\xi}_{i}^{(\mu)}(k)$ and $\tilde{\xi}_{l}^{(\mu)}(k)$? Here we use the conditional noise distribution in the natural dynamics of the system in order to keep the reference distribution close to optimal. Given the multivariate Gaussian distribution $\mathbb{G}(\xi_{i},\xi_{j},\xi_{l})$ for the original system of noises, and given that we have specified $\xi_{j}=\tilde{\xi}_{j}^{(\nu\mu)}(k)$, the distribution we need to sample from is given by the conditional distribution $\mathbb{G}(\xi_{i}=\tilde{\xi}_{i},\xi_{l}=\tilde{\xi}_{l}|\xi_{j}=\tilde{\xi}_{j}^{(\nu\mu)}$. Using the Schur complement formula, the mean and the covariance matrix of this distribution are given by $\Sigma_{0}\Sigma^{-1}\tilde{\xi}_{j}^{(\nu\mu)}$ and $2\widetilde{D}-\Sigma_{0}\Sigma^{-1}\Sigma_{0}^{T}$ respectively, where
$\Sigma_{0}$ $=2\begin{pmatrix}D_{ij}\\ D_{lj}\end{pmatrix},$ (S9)
$\Sigma$ $=2D_{jj},\mathrm{~{}and}$ (S10)
$\widetilde{D}$ $=\begin{pmatrix}D_{ii}&&D_{il}\\ D_{il}&&D_{ll}\end{pmatrix}.$ (S11)
Summarizing, the procedure for generating $X_{i}^{(\mu)}$ and $X_{l}^{(\mu)}$ trajectories for a given $X_{j}^{(\nu)}$ trajectory consists of the following steps at every timestep:
  • The drifts for propagating $X_{i}^{(\mu)}(k)$ and $X_{l}^{(\mu)}(k)$ are obtained using $X_{i}^{(\mu)}(k)$, $X_{j}^{(\nu)}(k)$ and $X_{l}^{(\mu)}(k)$.
  • From the difference between $X_{j}^{(\nu)}(k)$ and $X_{j}^{(\nu)}(k+1)$, $\dot{X}_{j}^{(\nu)}(k)$ is obtained.
  • From $X_{i}^{(\mu)}(k)$ and $X_{l}^{(\mu)}(k)$, together with $X_{j}^{(\nu)}(k)$, the drift in $\dot{X}_{j}^{(\nu)}(k)$ is obtained.
  • From $\dot{X}_{j}^{(\nu)}(k)$ and the drift in $X_{j}$, the noise $\tilde{\xi}_{j}^{(\nu\mu)}(k)$ is obtained.
  • With this noise $\tilde{\xi}_{j}^{(\nu\mu)}(k)$ specified, we can sample the noise $\tilde{\xi}_{i}(k)$ and $\tilde{\xi}_{l}(k)$.
  • Using the drifts and the noise, we propagate $X_{i}^{(\mu)}(k)$ and $X_{l}^{(\mu)}(k)$.
Although $X_{i}^{(\mu)}$ and $X_{l}^{(\mu)}$ trajectories become overall a complicated nonlinear function of the $X_{j}^{(\nu)}$ trajectory due to the conditional noise sampling, the computations in each step are linear and simple. For accessing conditional distributions of trajectories of other variables, the reference dynamics is worked out similarly. Thus for every marginal probability computation for the $\nu$-th trajectory, a unique reference dynamics is used, which is fine-tuned to that trajectory. This method of choosing a reference dynamics is a numerical analogue of constructing an approximation for the solution to the stochastic filtering equation3, albeit one whose error can be exactly corrected through trajectory reweighting (see also [4] for an exact solution to the filtering problem in a class of Poisson-type channels). This drastically reduces the computational cost and makes TE-PWS feasible and accurate.
Appendix E Computing transfer entropy for jump processes
Here we describe in full detail the computation of transfer entropy in jump processes. For clarity, we reproduce Eqs. 9-11 from main text here,
[…]
$\Delta\mathcal{T}_{[k,k+1]}^{i\to j}$ $=\left\langle\pi_{X_{i}\to X_{j}}-\pi_{X_{j}}\right\rangle$ (S14)
where $\pi_{X_{i}\to X_{j}}$ is a functional of the trajectory segments $X_{i,[k,k+1]}$ and $X_{j,[k,k+1]}$, $\pi_{X_{j}}$ is a functional of only the segment $X_{j,[k,k+1]}$, and $\Delta\mathcal{T}_{[k,k+1]}^{i\to j}$ is the transfer entropy from $X_{i}$ to $X_{j}$ over the duration from $k\delta t$ to $(k+1)\delta t$. Here $\alpha$ counts the jumps that change the state of $X_{j}$, $N_{j}$ in number, $\mathcal{Q}_{ij},\lambda_{ij}$ and $\mathcal{Q}_{j},\lambda_{j}$ are jump and escape propensities for $X_{j}$ in the marginal spaces of $(X_{i},X_{j})$ and $(X_{j})$ respectively. The marginal jump propensities $\mathcal{Q}_{ij}(\alpha)$ and $\mathcal{Q}_{j}(\alpha)$ in Eqs. S12 and S13 are abbreviations for the marginal jump propensities $\mathcal{Q}_{ij}(\alpha,t_{\alpha})$ and $\mathcal{Q}_{j}(\alpha,t_{\alpha})$ with $t_{\alpha}$ denoting the time of the $\alpha$-th jump along the $X_{j}$ trajectory. The marginal jump propensities for an arbitrary jump labeled as $\beta$ at an arbitrary time $t$ are defined as
$\mathcal{Q}_{ij}(\beta,t)$ $\equiv\int D[X_{l,[0,t]}]\;Q_{\beta}(t)\;P(X_{l,[0,t]}|X_{i,[0,t]},X_{j,[0,t]})$ (S15)
$\mathcal{Q}_{j}(\beta,t)$ $\equiv\int\int D[X_{i,[0,t]}]\;D[X_{l,[0,t]}]\;Q_{\beta}(t)\;P(X_{i,[0,t]},X_{l,[0,t]}|X_{j,[0,t]})$ (S16)
where $\mathcal{Q}_{\beta}(t)$ is the jump propensity in the full $d$-dimensional space, and, with a slight abuse of notation, $X_{i,[0,t]}$ denotes the trajectory of $X_{i}$ from time 0 to $t$. The marginal escape propensities $\lambda_{ij}(t)$ and $\lambda_{j}(t)$ in Eqs. S12 and S13 are defined by summing marginal jump propensities $\mathcal{Q}_{ij}(\beta,t)$ and $\mathcal{Q}_{j}(\beta,t)$ respectively over all possible jumps $\beta$ that change the state of $X_{j}$,
${\lambda}_{ij}(t)$ $\equiv\sum_{\beta}{\mathcal{Q}}_{ij}(\beta,t)$ (S17)
${\lambda}_{j}(t)$ $\equiv\sum_{\beta}{\mathcal{Q}}_{j}(\beta,t)$ (S18)
Though the transfer entropy $\Delta\mathcal{T}_{[k,k+1]}^{i\to j}$ is formally defined using both escape and jump propensity terms in Eq. S14, only the latter terms contribute to the transfer entropy on average, as the escape propensity terms cancel on average, $\langle\lambda_{ij}\rangle=\langle\lambda_{j}\rangle$5. This is because for any $\beta$-th jump at any time $t$, $\langle\mathcal{Q}_{ij}(\beta,t)\rangle=\langle\mathcal{Q}_{j}(\beta,t)\rangle$ where the angular brackets denote an average over all trajectories, as can be seen from Eqs. S15 and S16. Thus only the marginal jump propensities at the specific jump times of $X_{j}$, $\mathcal{Q}_{ij}(\alpha)$ and $\mathcal{Q}_{j}(\alpha)$, are formally needed for computing the transfer entropy5. For computing $\mathcal{Q}_{ij}(\alpha)$ and $\mathcal{Q}_{j}(\alpha)$ by marginalization of the jump propensity in the full $d$-dimensional space, $\mathcal{Q}_{\alpha}(t_{\alpha})$, over conditional distributions, as defined in Eqs. S15 and S16, we need the conditional probabilities of the trajectories up to the time of each jump $t_{\alpha}$. This is indeed available on-the-fly from the trajectory weights in the RR scheme, as explained in the End Matter (EM) of the main text. The only difference from the case of the diffusive processes is that in the current case, trajectory weights are computed on-the-fly not only up to times $k\delta t$ and $(k+1)\delta t$, but also up to all intermediate times $t_{\alpha}$ that represent the jump times in $X_{j}$. These weights reflect the conditional distributions $P(X_{l,[0,t_{\alpha}]}|X_{i,[0,t_{\alpha}]},X_{j,[0,t_{\alpha}]})$ and $P(X_{i,[0,t_{\alpha}]},X_{l,[0,t_{\alpha}]}|X_{j,[0,t_{\alpha}]})$. We then compute $\mathcal{Q}_{ij}$ and $\mathcal{Q}_{j}$ in Eqs. S12 and S13 respectively for the $\alpha$-th jump as weighted averages over the trajectories,
$\mathcal{Q}_{ij}(\alpha)$ $=\int D[X_{l,[0,t_{\alpha}]}]\;Q_{\alpha}(t_{\alpha})\;P(X_{l,[0,t_{\alpha}]}|X_{i,[0,t_{\alpha}]},X_{j,[0,t_{\alpha}]})$ (S19)
$\mathcal{Q}_{j}(\alpha)$ $=\int\int D[X_{i,[0,t_{\alpha}]}]\;D[X_{l,[0,t_{\alpha}]}]\;Q_{\alpha}(t_{\alpha})\;P(X_{i,[0,t_{\alpha}]},X_{l,[0,t_{\alpha}]}|X_{j,[0,t_{\alpha}]})$ (S20)
where $Q_{\alpha}(t_{\alpha})$, the jump propensity in the full $d$-dimensional space, is analytically available. Thus, using only the jump propensity terms in Eqs. S12-S14, we obtain an exact estimate of the transfer entropy, which we call $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J})}$. This approach of the computation of the transfer entropy using only the jump propensities has been recently shown to have a faster convergence over an alternate time-discretized approach for neural spike train data6.
Appendix F Reduced variance estimator in jump processes
Surprisingly, we find that the statistical error in the estimate $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J})}$ can be further reduced by an order of magnitude by adding to it the integral of the escape terms in Eqs. S12 and S13, marginalized with a quadrature of $\delta t$. This reduction arises because fluctuations in the jump terms in Eqs. S12 and S13 are pathwise anti-correlated to those in the escape terms, even though the latter cancel on average, as explained below.
The improved estimate is obtained by marginalizing the escape terms in Eq. S12 and S13 over conditional distributions $P\left(X_{l,[0,k]}\big{|}X_{i,[0,k]},X_{j,[0,k]}\right)$ and $P\left(X_{i,[0,k]},X_{l,[0,k]}\big{|}X_{j,[0,k]}\right)$ respectively, rather than $P\left(X_{l,[0,t]}\big{|}X_{i,[0,t]},X_{j,[0,t]}\right)$ and $P\left(X_{i,[0,t]},X_{l,[0,t]}\big{|}X_{j,[0,t]}\right)$ at every instant of time. Such a quadrature for the marginalization is necessary as the integral in the first terms in Eqs. S12 and S13 cannot be evaluated in closed form. The improved estimate, based on both the jump propensities J and the escape propensities E, is thus computed as
[…]
where $\mathcal{T}^{(\mathrm{J})}_{X_{i}\to X_{j}}$ is the estimate from the previous section based on only the jump propensity terms in Eqs. S12 and S13. In Eq. S21, ${\lambda}_{ij}(t)$ and ${\lambda}_{j}(t)$ are computed as ${\lambda}_{ij}(t)\equiv\sum_{\beta}{\mathcal{Q}}_{ij}(\beta,t)$, ${\lambda}_{j}(t)\equiv\sum_{\beta}{\mathcal{Q}}_{j}(\beta,t)$ with $\mathcal{Q}_{ij}(\beta,t)$ and ${\mathcal{Q}}_{j}(\beta,t)$ computed from Eqs. S15 and S16, respectively, but with
${P}(X_{l,[0,t]}|X_{i,[0,t]},X_{j,[0,t]})$ $\simeq P(X_{l,[0,k]}|X_{i,[0,k]},X_{j,[0,k]})\;P_{0}(X_{l,[k,t]}|X_{l,[0,k]})$ (S22)
${P}(X_{i,[0,t]},X_{l,[0,t]}|X_{j,[0,t]})$ $\simeq P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]})\;P_{0}(X_{i,[k,t]},X_{l,[k,t]}|X_{i,[0,k]},X_{l,[0,k]})$ (S23)
Here $X_{i,[0,k]}$ and $X_{i,[k,t]}$ denote the $X_{i}$ trajectory between times 0 and $k\delta t$ and between times $k\delta t$ and $t$ respectively; $P_{0}$ is, as before, the reference distribution from which we generate the trajectories of the hidden variables. Hence, we see that the estimator $\mathcal{T}^{(\mathrm{J}+\mathrm{E})}_{X_{i}\to X_{j}}$ is computed from the trajectory weights at time $k\delta t$ as available from the RR scheme, rather than the weight at every instant of time $t$. We refer to this as the quadrature approximation for marginalization, which incurs an $\mathcal{O}(\delta t)$ error.
[…]
$\Delta\mathcal{T}_{[k,k+1]}^{i\to j}$ $=\Bigg{\langle}Q_{ij}\delta t\ln\frac{Q^{ij}}{Q^{*}}-(Q_{ij}-Q^{*})\delta t\Bigg{\rangle}$ (S24)
where the angular brackets now denote an average over $X_{i}$ trajectories. We see that the second term fluctuates around a mean of 0. By Taylor expanding the first term upto second order in $(Q_{ij}-Q^{*})$, we find
$\Delta\mathcal{T}_{[k,k+1]}^{i\to j}$ $=\Bigg{\langle}(Q_{ij}-Q^{*})\delta t+\frac{\delta t(Q_{ij}-Q^{*})^{2}}{2Q^{*}}+\mathcal{O}\big{(}(Q_{ij}-Q^{*})^{3}\big{)}-(Q_{ij}-Q^{*})\delta t\Bigg{\rangle}$ (S25)
$=\Bigg{\langle}\frac{\delta t(Q_{ij}-Q^{*})^{2}}{2Q^{*}}+\mathcal{O}\big{(}(Q_{ij}-Q^{*})^{3}\big{)}\Bigg{\rangle}$ (S26)
So fluctuations in the second term of Eq. S24 cancel a part of the fluctuations in the first term. This results in a smaller overall variance compared to the case where only the first term of Eq. S24 is used. The above proof also holds in the case of multiple kinds of jumps in $X_{j}$, such as increments and decrements in copy numbers, because the escape propensities $\lambda_{ij}$ and $\lambda_{j}$ are additive over the different kinds of jumps and the above proof works separately for each kind of jump.
We numerically demonstrate this effect in a chemical reaction network of two species $X$ and $Y$, consisting of reactions $\phi\rightarrow X,X\rightarrow\phi,X\rightarrow X+Y,Y\rightarrow\phi$, with rate constants $k_{1}=50$, $k_{-1}=1$, $k_{2}=10$ and $k_{-2}=10$ respectively. Plotted in Figs. S1a and b are the two different transfer entropy estimates and their errors as a function of increasing trajectory duration and increasing statistical averaging respectively. As there is no feedback, the transfer entropy is formally equal to the exact mutual information estimate from PWS1, against which we have compared our results. We find that though both estimates yield unbiased results, the reduced-variance estimate $\mathcal{T}_{X\to Y}^{(\mathrm{J}+\mathrm{E})}$ has an order of magnitude smaller error than the jump-based estimate $\mathcal{T}_{X\to Y}^{(\mathrm{J})}$, for the same computational cost. On the other hand, the $\mathcal{O}(\delta t)$ error in $\mathcal{T}_{X\to Y}^{(\mathrm{J}+\mathrm{E})}$ from the quadrature in marginalization is negligible. $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J}+\mathrm{E})}$ is thus a more accurate estimate of the transfer entropy than $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J})}$. We expect this theoretical result to be tested using experimental data in the future, such as using data from neural spike trains6.
Figure S1: Accuracy of transfer entropy estimates $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J})}$ and $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J}+\mathrm{E})}$ in jump processes. (a) Transfer entropy as a function of time from the jump-based estimate $\mathcal{T}_{X\to Y}^{(\mathrm{J})}$ (red circles), reduced-variance estimate $\mathcal{T}_{X\to Y}^{(\mathrm{J}+\mathrm{E})}$ (blue triangles) and ground truth estimate from PWS (black dashed line) in the chemical reaction network with two species $X$ and $Y$. (b) Error at trajectory duration $T=0.2/k_{-1}$, defined as the difference of each estimate from the PWS estimate, as a function of increasing Monte-Carlo averaging. Plotted are errors in the jump-based estimate $\mathcal{T}_{X\to Y}^{(\mathrm{J})}$ (orange squares) and the reduced-variance estimate $\mathcal{T}_{X\to Y}^{(\mathrm{J}+\mathrm{E})}$ (green crosses). Initial conditions for the simulations are $n_{X}=50$ and $n_{Y}=500$, where $n_{X}$ and $n_{Y}$ are the number of species of $X$ and $Y$ respectively. $M_{1}$ for (a) is 25600. $M_{2}=1000$ for both subfigures.
Appendix G Variants of transfer entropy
A central advance in TE-PWS is the computation of the probability of a dynamical fluctuation by marginalization over trajectories of hidden variables. Here we show, in the context of discretized Langevin processes, how the same approach can be used to compute other trajectory-based metrics of directional information transmission beyond Schreiber’s transfer entropy.
Directed information. An alternate measure of information transmission from the trajectory $X_{i,[0,N]}$ to $X_{j,[0,N]}$ is given by directed information 7, defined as
$I[X_{i,[0,N]}\to X_{j,[0,N]}]=\sum_{k=0}^{N-1}I\left(X_{j}(k+1);X_{i,[0,k+1]}|X_{j,[0,k]}\right)$ $=\Bigg{\langle}\ln\frac{\prod_{k}P\left(X_{j}(k+1)|X_{i,[0,k+1]},X_{j,[0,k]}\right)}{\prod_{k}P\left(X_{j}(k+1)|X_{j,[0,k]}\right)}\Bigg{\rangle}$ (S27)
By comparing the above definition with that of the transfer entropy in Eq. 2 of the main text, we see that the directed information incorporates the stepwise information transmission into $X_{j}$ coming from the entire trajectory of $X_{i}$ including the current value $X_{i}(k+1)$, while Schreiber’s transfer entropy excludes the flow from the current value of $X_{i}$ (compare Eq. S27 against Eqs. 2 and 4 in the main text). We can bring the directed information to a computable form by taking $X_{i}(k+1)$ out of the conditioning in the numerator,
$I[X_{i,[0,N]}\to X_{j,[0,N]}]$ $=\Bigg{\langle}\ln\frac{\prod_{k}P\left(X_{j}(k+1)|X_{i,[0,k+1]},X_{j,[0,k]}\right)}{\prod_{k}P\left(X_{j}(k+1)|X_{j,[0,k]}\right)}\Bigg{\rangle}$
$=\Bigg{\langle}\ln\prod_{k}P\left(X_{i}(k+1),X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]}\right)\Bigg{\rangle}-\Bigg{\langle}\ln\prod_{k}P\left(X_{j}(k+1)|X_{j,[0,k]}\right)\Bigg{\rangle}$
$\qquad\qquad-\Bigg{\langle}\ln\prod_{k}P\left(X_{i}(k+1)|X_{i,[0,k]},X_{j,[0,k]}\right)\Bigg{\rangle}$ (S28)
Here we note that only two marginalization integrals are actually required for computing the three probabilities in Eq. S28. One is over the conditional distribution $P(X_{l,[0,k]}|X_{i,[0,k]},X_{j,[0,k]})$ and the other over $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]})$, exactly the same as those sampled in TE-PWS for calculating $\mathcal{T}_{X_{i}\to X_{j}}$, where $X_{l}$ denote all variables except $X_{i}$ and $X_{j}$. These two distributions can give the probabilities in Eq. S28 as
$P\left(X_{i}(k+1),X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]}\right)$ $=\int D[X_{l,[0,k]}]\;P\left(X_{l,[0,k]}|X_{i,[0,k]},X_{j,[0,k]}\right)$
.$\qquad.\;P\left(X_{i}(k+1),X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]},X_{l,[0,k]}\right)$ (S29)
$P\left(X_{i}(k+1)|X_{i,[0,k]},X_{j,[0,k]}\right)$ $=\int D[X_{l,[0,k]}]\;P\left(X_{l,[0,k]}|X_{i,[0,k]},X_{j,[0,k]}\right)$
.$\qquad.\;P\left(X_{i}(k+1)|X_{i,[0,k]},X_{j,[0,k]},X_{l,[0,k]}\right)$ (S30)
$P\left(X_{j}(k+1)|X_{j,[0,k]}\right)$ $=\int\int D[X_{i,[0,k]}]\;D[X_{l,[0,k]}]\;P\left(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]}\right)$
.$\qquad.\;P\left(X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]},X_{l,[0,k]}\right)$ (S31)
where, aside from the two conditional distributions $P(X_{l,[0,k]}|X_{i,[0,k]},X_{j,[0,k]})$ and $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]})$, all the other probabilities in the integrals are analytically available. Thus, by sampling the two conditional distributions through the RR scheme, computing the three averages in Eqs. S29-S31 as averages over those conditional distributions, plugging them in Eq. S28, and evaluating Eq. S28 as a single Monte-Carlo average, TE-PWS can compute directed information with the same computational cost as transfer entropy.
Conditional transfer entropy. The conventional transfer entropy can have a positive value even when there is no direct causal link from the input to the output variable, when information is being causally transmitted through intermediate variables. This motivated the definition of a conditional transfer entropy, also known as causation entropy, that can measure direct causal links8, 9, 10. For any choice of a third variable $X_{m}$, the conditional transfer entropy is defined as
$\mathcal{T}_{X_{i}\to X_{j}|X_{m}}=\sum_{k=0}^{N-1}I\left(X_{j}(k+1);X_{i,[0,k]}|X_{j,[0,k]},X_{m,[0,k]}\right)=\sum_{k}\Bigg{\langle}\ln\frac{P\left(X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]},X_{m,[0,k]}\right)}{P\left(X_{j}(k+1)|X_{j,[0,k]},X_{m,[0,k]}\right)}\Bigg{\rangle}$ (S32)
This expectation can be computed similar to the ordinary transfer entropy in Eq. 4 of the main text. The average is computed in a Monte-Carlo fashion over simulated trajectories of all variables. For each set of trajectories, the numerator and denominator are computed by marginalizing over all other hidden variables $X_{l}$ which exclude $X_{m}$ this time. The optimal reference dynamics should now be chosen to incorporate the effects of the $X_{m}$ trajectory through a frozen field of drift and diffusion resulting from the fixed $X_{m}$ trajectory, similar to how the $X_{i}$ and $X_{j}$ trajectories influence the reference dynamics as discussed above, under Choice of reference distribution. Thus, calculation of each conditional transfer entropy with TE-PWS requires two marginalization integrals, similar to the ordinary transfer entropy.
Filtered transfer entropy. Recently, filtered transfer entropy has been proposed as a way to quantify information transfer in the spirit of filtering theory11. The filtered transfer entropy from $X_{i}$ to $X_{j}$ is defined as
$\widehat{\mathcal{T}}_{X_{i}\to X_{j}}=\sum_{k=0}^{N-1}I\left(X_{i}(k+1);X_{j}(k+1)|X_{j,[0,k]}\right)$
$=\sum_{k}\Bigg{\langle}\ln\frac{P\left(X_{i}(k+1),X_{j}(k+1)|X_{j,[0,k]}\right)}{P\left(X_{i}(k+1)|X_{j,[0,k]}\right)P\left(X_{j}(k+1)|X_{j,[0,k]}\right)}\Bigg{\rangle}$ (S33)
which quantifies how much the prediction of $X_{i}(k+1)$ is improved by using $X_{j}(k+1)$ in addition to the past trajectory $X_{j,[0,k]}$. The computation of $\widehat{\mathcal{T}}_{X_{i}\to X_{j}}$ requires marginalization over only one conditional distribution, $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]})$. Each of the probabilities in Eq. S33 can be computed by averaging analytically available transition probabilities $P\left(X_{i}(k+1),X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]},X_{l,[0,k]}\right)$, $P\left(X_{i}(k+1)|X_{i,[0,k]},X_{j,[0,k]},X_{l,[0,k]}\right)$ and $P\left(X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]},X_{l,[0,k]}\right)$, over this conditional distribution $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]})$, which is provided by TE-PWS through the RR scheme. Thus, TE-PWS can be used to compute filtered transfer entropy at half the computational cost as Schreiber’s transfer entropy.